OpenStudy (anonymous):

PLEASE will give medal!! Simplify sin Θ/square root of 1-sin^2 Θ

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Answer choices Tan^2 Θ -1 Tan Θ 1

OpenStudy (anonymous):

@kingofgames

Nnesha (nnesha):

\[\huge\rm \frac{ \sin \theta }{ \sqrt{ 1 -\sin^2 \theta} }\] 1 -sin^2 = ??

OpenStudy (anonymous):

Thats what it looks like. Thanks I have no idea

Nnesha (nnesha):

\[\huge\rm sin^2 \theta + \cos^2 \theta =1\] trig identity can you solve it for cos^2 ?

OpenStudy (anonymous):

So the bottoms equals 0?

Nnesha (nnesha):

nope \[\huge\rm sin^2 \theta + \cos^2 \theta =1\] subtract sin^2 both side

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

well theres 1- then the 1 that you said we get above which equals 0... or nah

Nnesha (nnesha):

nope let's say sin = x and cos = y  how would you solve this equation for y ?